OpenStudy (anonymous):

Part 1: Solve the following equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 10x + 5 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (radar):

I choose "completing the square" method because it is not factorable, and the coefficient of x is a 1. x^2 +10x +25=-5+25 (x+5)^2=20 \[x+5=\sqrt{20}=2\sqrt{5}\] \[x=-5\pm2\sqrt{5}\]

OpenStudy (radar):

typo I meant to say the coefficient of x^2 is a 1.